A person kicks a ball with an initial velocity 15m/s at an angle 37° above the horizontal (neglect

the air resistance). Find:

a. the total time the ball is in the air.

b. the horizontal distance traveled by the ball.


1
Expert's answer
2021-11-09T10:59:05-0500

a) The total time is given as follows (see https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/):


"T = \dfrac{2u\sin\theta}{g}"

where "u = 15m/s, \theta = 37\degree, g = 9.8m/s^2". Thus, obtain:


"T = \dfrac{2\cdot 15\cdot \sin37\degree}{9.8} \approx 1.8s"

b) The horizontal distance:


"R = \dfrac{u^2\sin2\theta}{g} = \dfrac{15^2\cdot \sin(2\cdot 37\degree)}{9.8} \approx 22m"

Answer. a) 1.8s, b) 22m.